Nesmith Corporation's outstanding bonds have a $1,000 par value, a 8% semiannual coupon, 14 years to maturity, and an 7% YTM. What is the bond's price? Round your answer to the nearest cent.
Answer #1

Information provided:

Face value= future value= $1,000

Time= 14 years*2= 28 semi-annual periods

Coupon rate= 8%/2= 4% per semi-annual period

Coupon payment= 0.04*1,000= $40

Yield to maturity= 7%/2= 3.50% per semi-annual period

The price of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

N= 28

PMT= 40

I/Y= 3.50

Press the CPT key and PV to compute the present value.

The value obtained is 1,088.34.

Therefore, the market price of the bond is $1,088.34.

In case of any query, kindly comment on the solution.